Project

General

Profile

NDOS CommissionMtng 20120806

Attending: Gavin, Kurt, Mat, Leon, Jan, Susan and Nick

UPDATES:
1.  Detector Update
Msg log disk filled last week: Gennadiy had made some changes to the alarming that may have lead to this. (Kurt will ask for some more details.)
T-Storm took run down on Saturday

2.  DataCheck Update
Jan- July 27-Aug 4 (reprocessing stuck there after)

FEB with repeated shut 2-1:21 3-2:23

Tuesday afternoon - Data disk filled: Jan saw data flow stop to the DataCheck.
Friday to Monday

FULL REPORT:
Run 14010
From: 2012-07-27 17:28:38
To: 2012-07-28 17:26:38
24 Subruns
FEB3-2:27 shutoff in subrun 15

Run 14011
From: 2012-07-28 17:26:38
To: 2012-07-29 17:24:39
24 Subruns
FEB3-1:8 shutoff in subrun 6
FEB3-2:13 shutoff in subrun 20

Run 14012
From: 2012-07-29 17:24:39
To: 2012-07-30 11:19:25
18 Subruns
FEB2-1:21 shutoff in subrun 6
FEB3-2:23 shutoff in subrun 8

Monday to Wednesday
Run 14013
From: 2012-07-30 11:56:12
To: 2012-07-31 11:54:11
24 Subruns
FEB3-2:23 shutoff in subrun 0
FEB2-1:21 shutoff in subrun 4
FEB3-1:8 shutoff in subrun 8

Additionally Jan reported:
14021 go belly up at 2012-08-04 16:02:11, presumably due to the thunderstorms. The problem is, it was not caught by my, now automated, script because it was not an overnight run. The overnight run 14020 rolled over into 14021, I'll add a few lines to look for those types of runs.

4.  Installation Plans for the week
Leon requests - some failed APDs be removed 3-6
Some APD terminators have been repaired at Fermilab and will be tested this week
Hamatsu- They don't think our problem is related to their cycling issue.
How much work is there for Ron to get a version? Kurt indicated about 1-2 weeks. Leon agree that this was an appropriate time scale.

5.  DAQ Update and Plans for the week
AppMan changes, DCM startup scripts, DCS Cooling speed, Gennadiy and Steven test the new email and dryer system alarms. Tuesday and Wednesday and possibly Thursday.

6.  DCS Update and Plans for the week
See Kurt.

7.  AOB